# Entropic Risk Optimization in Discounted MDPs: Sample Complexity Bounds with a Generative Model

### v2

@@ 12    $\forall s\in\mathcal{S}:\widehat{P}(s,z)=\frac{m(s,z)}{n}$
The model-based approach we describe is general in the sense that any oracle that for any $\varepsilon>0$ can find an $\varepsilon$-optimal policy can be used. We prove the existence of one such oracle in the form of a $Q$-value iteration very like the one from the classical risk-neutral setting. The proof is very similar to Part (a) in Theorem 3.1 in asienkiewicz2017note but is nevertheless provided in Appendix F.
Lemma 1 (Q-value iteration).
Fix a map $\pi:\mathcal{A}\rightarrow\mathcal{S}$. We then define the operators $\mathcal{T}^{\pi},\mathcal{T}:\mathbb{R}^{S\times A}\rightarrow\mathbb{R}^{S\times A}$ which for $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is given by
$(\mathcal{T}f)(s,a)$ $=R(s,a)+\frac{-\gamma}{\beta}\log\bigg(\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta\max_{a}f(s^{\prime},a)}\bigg),$
$(\mathcal{T}^{\pi}f)(s,a)$ $=R(s,a)+\frac{-\gamma}{\beta}\log\bigg(\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta f(s^{\prime},\pi(s^{\prime}))}\bigg).$
The operators $\mathcal{T}$ and $\mathcal{T}^{\pi}$ are $\gamma$-contractions with respect to the max-norm, i.e., for value-functions $f_{1}$ and $f_{2}$, it holds that $\|\mathcal{T}f_{1}-\mathcal{T}f_{2}\|\leq\gamma\|f_{1}-f_{2}\|$ and $\|\mathcal{T}^{\pi}f_{1}-\mathcal{T}^{\pi}f_{2}\|\leq\gamma\|f_{1}-f_{2}\|$.
The above lemma is the basis for the Q-value iteration algorithm:
@@
Input: MDP $M=(\mathcal{S},\mathcal{A},P,R,\gamma,\beta)$ and number of iterations $k$.
@@ 2 Initialization: $\forall(s,a)$ set $Q(s,a)=0$
3 for j=0,1,,k1j=0,1,...,k-1 do
4  for all $s\in\mathcal{S}$ do
5     $\pi_{j}(s)=\arg\!\max_{a\in\mathcal{A}}Q_{j}(s,a)$
6     for all $a\in\mathcal{A}$ do
7        $\mathcal{T}Q_{j}(s,a)=R(s,a)-\frac{\gamma}{\beta}\log(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta Q_{j}(s,\pi_{j}(s))}])$
8        $Q_{j+1}(s,a)=\mathcal{T}Q_{j}(s,a)$
@@
Algorithm 2 RS-QVI(M,k)(M,k)
The next lemma shows that if we choose $k$ large enough in the RS-QVI algorithm, we can obtain $Q_{k}$ and $V^{\pi_{k}}$ that are as close to $Q^{*}$ and $V^{*}$ as we desire. The proof is postponed to Appendix F.
@@
Fix $\varepsilon>0$. Then there exists some $k(\varepsilon)$ such that if the number of iterations in RS-QVI exceeds $k(\varepsilon)$, then the output of Algorithm 2 (RS-QVI) satisfies $\|Q_{k}-Q^{*}\|<\varepsilon$ and $\|V^{\pi_{k}}-V^{*}\|<\varepsilon$.
Using Algorithm 2, we introduce the MB-RS-QVI algorithm, which consists in building an empirical model $\widehat{M}=(\mathcal{S},\mathcal{A},\widehat{P},R,\gamma,\beta)$ via calling the generative model $n$ times – namely, $\widehat{P}=\texttt{EstimateModel}(n)$ – and then solving it via RS-QVI.
3.1 Analysis of MB-RS-QVI
With RS-QVI in place, we need a set of lemmas for the analysis of the sample-complexity of the model based RS-QVI algorithm.
An important result for the analysis is a risk-sensitive version of the simulation lemma kearns2002near ; strehl_analysis_2008 , which describes how different two $Q$-functions for the same policy are in two different MDPs that differ only slightly in their rewards and transition functions. The proof is postponed to Appendix F.
Lemma 3 (Simulation Lemma with Entropic Risk).
Consider two MDPs $M_{1}=(S,A,P_{1},R,\gamma,\beta)$ and $M_{2}=(S,A,P_{2},R,\gamma,\beta)$ differing only in their transition functions. Fix a stationary policy $\pi$, and let $Q^{\pi}_{1}$ and $Q^{\pi}_{2}$ be respective $Q$-functions of $\pi$ in $M_{1}$ and $M_{2}$. If $\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\|P_{1}(\cdot|s,a)-P_{2}(\cdot|s,a)\|_{1}\leq\tau$. Then, it holds that
$\|Q^{\pi}_{1}-Q^{\pi}_{2}\|\leq\begin{cases}\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-|\beta|(\frac{1}{1-\gamma}-V_{1}(s^{\prime}))}\big|\quad&\beta<0\\ \frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-|\beta|V_{1}(s^{\prime}))}\big|\quad&\beta>0\\ \frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\max_{s,a}\sum_{s^{\prime}\in\mathcal{S}}|P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)|\quad&\beta\neq 0.\end{cases}$
It then follows that if for some $\varepsilon>0$ it holds that any of the $\max_{s,a}$-expressions is smaller than $\varepsilon\frac{1-\gamma}{\gamma}|\beta|e^{-|\beta|\frac{1}{1-\gamma}}$, then $\|Q_{1}^{\pi}-Q_{2}^{\pi}\|\leq\varepsilon$ and ensuring this with high probability is a matter of invoking an appropriate concentration inequality. Using the decompositions from Lemma 4, we get for any $s\in\mathcal{S}$ that
$V^{\pi_{k}}(s)$ $\geq V^{*}(s)-\|V^{\pi_{k}}-\widehat{V}^{\pi_{k}}\|-\|\widehat{V}^{\pi^{*}}-V^{*}\|-\|\widehat{V}^{\pi_{k}}-\widehat{V}^{*}\|,$ (10)
and similarly we get for any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ that
$Q_{k}(s,a)\geq Q^{*}(s,a)-\|\widehat{Q}^{\pi^{*}}-Q^{*}\|-\|Q_{k}-\widehat{Q}^{*}\|\,.$ (11)
The last of the distances on the right-hand side of (10) and (11) can be made arbitrarily small by any optimization oracle for the problem. One such is value-iteration using the model MDP as demonstrated by Lemma 2. Making these terms small enough is thus purely a computational matter and not a statistical one. We thus focus on bounding the remaining distances on the right-hand sides.
Theorem 1.
There exists a universal constant $c$ such that for any $\varepsilon>0,\delta\in(0,1)$ and any MDP $M$ with $S$ states and $A$ actions, if the learner is allowed to make
$T=c\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|}\bigg)^{2}\log\bigg(\frac{SA}{\delta}\bigg)$ (12)
@@ 3.1 Analysis of MB-RS-QVI
calls to the generative model, then $\mathbb{P}(\|Q^{*}-Q_{k}\|\leq\varepsilon)\geq 1-\delta$.
@@
For any $\varepsilon>0$, we can get $\|Q_{k}-\widehat{Q}^{*}\|<\varepsilon/2$ using enough iterations of the optimization oracle by Lemma 2. The term $\|\widehat{Q}^{\pi^{*}}-Q^{*}\|$ can also be made smaller than $\varepsilon/2$ by the simulation lemma if either $\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-|\beta|(\frac{1}{1-\gamma}-V_{1}(s^{\prime}))}\big|<\tau$ or $\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-|\beta|V_{1}(s^{\prime}))}\big|<\tau$, where $\tau=\frac{\varepsilon}{2}\bigg[\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\bigg]^{-1}$ which can be ensured with probability larger than $1-\delta$ by picking $N=\frac{(1-e^{-\beta\frac{1}{1-\gamma}})^{2}}{2\tau^{2}}\log(2SA/\delta)$. Using that the total calls to the generative model is $T=SAN$ and substituting in the value for $\tau$, we can ensure for all $(s,a)$ that $Q_{k}(s,a)>Q^{*}(s,a)-\varepsilon$ with probability larger than $1-\delta$ by using a total number of samples
$T=3\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|}\bigg)^{2}\log\bigg(\frac{SA}{\delta}\bigg)\,.$ (13)
Theorem 2.
There exists a universal constant $c$ such that for any $\varepsilon>0,\delta\in(0,1)$, and any MDP $M$ with $S$ states and $A$ actions if the learner is allowed to make
$T=c\frac{SA}{\varepsilon^{2}(1-\gamma)^{4}}\bigg(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|}\bigg)^{2}\log\bigg(\frac{SA}{\delta}\bigg)$ (14)
calls to the generative model then $\mathbb{P}(\|V^{*}-V^{\pi_{k}}\|\leq\varepsilon)\geq 1-\delta$

### v3

@@ 13    $\forall s\in\mathcal{S}:\widehat{P}(s,z)=\frac{m(s,z)}{N}$
Input: Empirical MDP $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widehat{M}=(\mathcal{S},\mathcal{A},\widehat{P},R,\gamma,\beta)$}, number of iterations $k$
@@ 2 Initialization: $\forall(s,a)$ set $Q(s,a)=0$
3 for $j=0,1,\ldots,k-1$ do
4  for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ do
5     $Q_{j+1}(s,a)=R(s,a)-\frac{\gamma}{\beta}\log\big(\mathbb{E}_{s^{\prime}\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widehat{P}_{s,a}}}\big[e^{-\beta\max_{a^{\prime}}Q_{j}(s^{\prime},a^{\prime})}\big]\big)$
6
@@
Algorithm 2 RS-QVI
Let $\widehat{P}$ denote the plug-in estimator built using the $T=NSA$ independent samples obtained from the generative model; that is, for $(s,a,s^{\prime})\in Z\times\mathcal{S}$, $\widehat{P}(s^{\prime}|s,a)=\frac{n(s,a,s^{\prime})}{N}$, where $n(s,a,s^{\prime})$ denotes the number of times $s^{\prime}$ was observed under the queried pair $(s,a)\in Z$. The model-based approach we describe relies on the empirical MDP formulated using $\widehat{P}$, $\widehat{M}=(\mathcal{S},\mathcal{A},R,\widehat{P},\gamma,\beta)$, but it is otherwise general in the sense that it can use any oracle that outputs an $\varepsilon$-optimal policy for any $\varepsilon>0$. We prove the existence of one such oracle in the analysis (cf. Lemma 2) in the form of a $Q$-value iteration akin to that of the classical risk-neutral setting. It is the basis for RS-QVI (Algorithm 2), which is a $Q$-value iteration for discounted MDPs with ERM with correctness guarantees.
Equipped with these, we introduce MB-RS-QVI. For an input $\varepsilon>0$, the algorithm consists in:
  • (i)
    building an empirical MDP $\widehat{M}=(\mathcal{S},\mathcal{A},\widehat{P},R,\gamma,\beta)$ via calling the generative model $N$ times (namely, $\widehat{P}=\texttt{EstimateModel}(N)$);
  • (ii)
    returning RS-QVI(M^,k)($\widehat${M},k).
We show in Lemma 2 how to pick $k:=k(\varepsilon)$ to ensure $\varepsilon$-value-correctness (i.e., $\|Q^{*}-Q_{k}\|\leq\varepsilon$) and $\varepsilon$-policy-correctness (i.e., $\|V^{*}-V^{\pi_{k}}\|\leq\varepsilon$).
5 Sample Complexity Analysis of MB-RS-QVI
In this section, we present a sample complexity analysis of MB-RS-QVI under both policy learning and value learning.
5.1 Properties of RS-QVI
We first state two results for RS-QVI, establishing its convergence properties. Their proofs are presented in Appendix D.
Lemma 1 (Contraction properties).
The operators $\mathcal{T}$ and $\mathcal{T}^{\pi}$ are $\gamma$-contractions with respect to the max-norm, i.e., $\|\mathcal{T}f_{1}-\mathcal{T}f_{2}\|\leq\gamma\|f_{1}-f_{2}\|$ and $\|\mathcal{T}^{\pi}f_{1}-\mathcal{T}^{\pi}f_{2}\|\leq\gamma\|f_{1}-f_{2}\|$ for value functions $f_{1}$ and $f_{2}$.
The proof of this result is very similar to that of Part (a) in Theorem 3.1 in [4]; nevertheless, we include it for completeness. The next lemma shows that for large enough $k$ in RS-QVI, we can obtain $Q_{k}$ and $V^{\pi_{k}}$ that are as close to, respectively, $Q^{*}$ and $V^{*}$ as desired:
@@ 5.1 Properties of RS-QVI
Fix $\varepsilon>0$. Under RS-QVI (Algorithm 2), one has: (i) $\|Q_{k}-Q^{*}\|<\varepsilon$ if $k\geq\frac{-\log((1-\gamma)\varepsilon)}{\log(1/\gamma)}$; (ii) $\|V^{\pi_{k}}-V^{*}\|<\varepsilon$ if $k\geq\frac{\log 2-\log((1-\gamma)^{2}\varepsilon)}{\log(1/\gamma)}$.
5.2 Sample Complexity Upper Bounds
We are ready to present sample complexity bounds for MB-RS-QVI: Theorem 1 states such a result for the case of value learning, while Theorem 2 offers a bound for policy learning.
Theorem 1 (Sample complexity, value learning).
For any $\varepsilon>0$, $\delta\in(0,1)$, and any MDP $M$ with $S$ states and $A$ actions, if the learner makes
$T\geq\frac{2SA\gamma^{2}}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(\frac{e^{|\beta|/(1-\gamma)}-1}{|\beta|}\bigg)^{2}\log\bigg(\frac{SA}{\delta}\bigg)$
model calls to the generative model, then $\mathbb{P}(\|Q^{*}-Q_{k}\|\leq\varepsilon)\geq 1-\delta$.
Theorem 2 (Sample complexity, policy learning).
For any $\varepsilon>0,\delta\in(0,1)$, and any MDP $M$ with $S$ states and $A$ actions, if the learner makes
$T\geq\frac{9SA\gamma^{2}}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(\frac{e^{|\beta|/(1-\gamma)}-1}{|\beta|}\bigg)^{2}\min\Big\{\frac{\gamma^{2}}{(1-\gamma)^{2}}\log\bigg(\frac{4SA}{\delta}\bigg),\,\log\bigg(\frac{4SA|\Pi|}{\delta}\bigg)\Big\}$
@@ 5.2 Sample Complexity Upper Bounds
model calls, then $\mathbb{P}(\|V^{*}-V^{\pi_{k}}\|\leq\varepsilon)\geq 1-\delta$. Here, $\Pi$ denotes the set of stationary deterministic policies.
The sample complexity bound offered by Theorem 2 can be further simplified using the worst-case bound $|\Pi|\leq A^{S}$ to $\widetilde{\cal O}\big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\min\big\{S,\frac{1}{(1-\gamma)^{2}}\big\}\big)L^{2}\big)$ with $L=\frac{1}{|\beta|}\big(e^{|\beta|/(1-\gamma)}-1\big)$. To be more precise, by including log-terms, depending on whether $S\ll 1/(1-\gamma)^{2}$ in the problem at hand, one may obtain a bound of
${\cal O}\bigg(\frac{SA}{\varepsilon^{2}(1-\gamma)^{4}}\log\bigg(\frac{4SA}{\delta}\bigg)L^{2}\bigg)\qquad\hbox{or}\qquad{\cal O}\bigg(\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(S+\log\bigg(\frac{4SA}{\delta}\bigg)\bigg)L^{2}\bigg)\,.$
Let us remark however that in problems where $|\Pi|$ grows polynomially with $S$, one will get a substantially better bound.
Remark 2.
Taking the limit in the PAC bound of Theorems 12, as $\beta\rightarrow 0$, yields corresponding sample complexity bounds for the risk-neutral case. The resulting bound for value learning is off the optimal bound by a factor of $(1-\gamma)^{-1}$, and for policy learning by a factor of $(1-\gamma)^{-1}\min\{S,(1-\gamma)^{-2}\}$; we again refer to Table 1 for a related comparison. It is worth stressing, however, that these implied bounds are valid for the entire $\varepsilon$-range, unlike the results in, e.g., [30, 1, 67].
Remark 3.
Existing derivations of minimax sample complexity bounds in the risk-neutral setting (e.g., [30, 1, 67, 49]) rely on techniques that crucially exploit the additive structure of the return with respect to rewards, such as variance lemmas establishing Bellman consistency of the variance of cumulative discounted rewards. These tools do not extend to ERM due to its intrinsic non-linearity, and are therefore not applicable in our setting.
6 Proofs: Sample Complexity Upper Bounds
In this section, we prove Theorems 1 and 2. As preliminarily, we present some results that will be used in the proofs. The first one concerns basic decompositions of the error terms associated to $V^{\pi_{k}}$ and $Q_{k}$. Let $\widehat{V}^{*}$ and $\widehat{Q}^{*}$ denote the optimal value and Q-value in $\widehat{M}$, respectively, and $\pi^{*}$ denote an optimal policy in $M$. Then, for any $(s,a)\in\mathcal{S}\times\mathcal{A}$,
$Q_{k}(s,a)$ $\geq Q^{*}(s,a)-\|\widehat{Q}^{\pi^{*}}-Q^{*}\|-\|Q_{k}-\widehat{Q}^{*}\|\,$, (8)
$V^{\pi_{k}}(s)$ $\geq V^{*}(s)-\|V^{\pi_{k}}-\widehat{V}^{\pi_{k}}\|-\|\widehat{V}^{\pi^{*}}-V^{*}\|-\|\widehat{V}^{\pi_{k}}-\widehat{V}^{*}\|\,.$ (9)
These follow from standard techniques, but for completeness, we derive them in Lemma 4 in Appendix C. In (8), the term $\|\widehat{Q}^{\pi^{*}}\!-Q^{*}\|$ captures the statistical hardness due to having the generative model, whereas the term $\|Q_{k}-\widehat{Q}^{*}\|$ represents the computational challenge and can be made desirably small after enough iterations of value iteration, and its control follows from the contraction property of ERM, which is also present in the case of known model. Similarly, in (9), the terms $\|V^{\pi_{k}}-\widehat{V}^{\pi_{k}}\|$ and $\|\widehat{V}^{\pi^{*}}-V^{*}\|$ correspond to the statistical hardness, while $\|\widehat{V}^{\pi_{k}}-\widehat{V}^{*}\|$ captures the computational hardness.
The second result concerns smoothness of Q-values under ERM when the transition function is perturbed. More specifically, it asserts how different $Q$-values of a fixed policy are in two different MDPs that differ only slightly in their transition functions. This parallels the result in [41, 70] to ERMs.
Lemma 3 ($Q$-value smoothness under ERM).
Consider two MDPs $M=(S,A,P,R,\gamma,\beta)$ and $\widetilde{M}=(S,A,\widetilde{P},R,\gamma,\beta)$ differing only in their transition functions. Fix a stationary policy $\pi$, and let $Q^{\pi}$ and $\widetilde{Q}^{\pi}$ be respective $Q$-values of $\pi$ in $M$ and $\widetilde{M}$. It holds that $\|Q^{\pi}-\widetilde{Q}^{\pi}\|\leq\xi W_{1}$ for $\beta<0$, and $\|Q^{\pi}-\widetilde{Q}^{\pi}\|\leq\xi W_{2}$ for $\beta>0$, where $\xi:=\frac{\gamma}{(1-\gamma)|\beta|}e^{|\beta|/(1-\gamma)}$ and
$W_{1}$ $:=\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{s,a}(s^{\prime})-\widetilde{P}_{s,a}(s^{\prime})]e^{-|\beta|(\frac{1}{1-\gamma}-V^{\pi}(s^{\prime}))}\big|,$
$W_{2}$ $:=\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{s,a}(s^{\prime})-\widetilde{P}_{s,a}(s^{\prime})]e^{-|\beta|V^{\pi}(s^{\prime})}\big|\,,\qquad\hbox{with $V^{\pi}(s)=\max_{a}Q^{\pi}(s,a)$}.$
6.1 Proof of Theorem 1
@@
Let $\beta>0$ and $\varepsilon>0$. In view of the error decomposition in (8), to establish $\varepsilon$-value-correctness it suffices to ensure $\|\widehat{Q}^{\pi^{*}}-Q^{*}\|\leq\varepsilon/2$ and $\|Q_{k}-\widehat{Q}^{*}\|\leq\varepsilon/2$. By Lemma 2, we can have $\|Q_{k}-\widehat{Q}^{*}\|<\varepsilon/2$ using enough iterations of the optimization oracle. Further, by Lemma 3, if
$\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{s,a}(s^{\prime})-\widehat{P}_{s,a}(s^{\prime})]e^{-|\beta|V^{*}(s^{\prime})}\big|<\frac{\varepsilon(1-\gamma)|\beta|}{2\gamma}e^{-|\beta|/(1-\gamma)}=:\tau,$ (10)
then $\|\widehat{Q}^{\pi^{*}}-Q^{*}\|\leq\varepsilon/2$. To ensure this, we use the following lemma (proven in Appendix D):
{restatable}
lemmaHoeffdingERM Let $\pi$ be any fixed policy and $\tau>0$. If $N>\frac{1}{2\tau^{2}}\big(1-e^{-|\beta|/(1-\gamma)}\big)^{2}\log(2SA/\delta)$, then it holds that
$(i)$ $\qquad\max_{s,a}\big|\sum_{s^{\prime}}[P_{s,a}(s^{\prime})-{\widehat{P}_{s,a}(s^{\prime})}]e^{-\beta V^{\pi}(s^{\prime})}\big|<\tau\,,\quad\hbox{with probability $\geq 1-\delta$,}\quad\hbox{$\beta>0$};$
$(ii)$ $\qquad\max_{s,a}\big|\sum_{s^{\prime}}[P_{s,a}(s^{\prime})-{\widehat{P}_{s,a}(s^{\prime})}]e^{-|\beta|(V^{\pi}(s^{\prime})-\frac{1}{1-\gamma})}\big|<\tau\,\quad\hbox{with probability $\geq 1-\delta$,}\quad\hbox{$\beta<0$}.$
Applying Lemma 6.1, inequality (i), with $\pi=\pi^{\star}$, we observe that (10) holds with probability at least $1-\delta$ by picking $N\geq\frac{1}{2\tau^{2}}\big(1-e^{-|\beta|/(1-\gamma)}\big)^{2}\log(2SA/\delta)$. Noting that the total calls to the generative model is $T=SAN$ and substituting in the value for $\tau$, we can ensure for all $(s,a)$ that $Q_{k}(s,a)>Q^{*}(s,a)-\varepsilon$ with probability larger than $1-\delta$ by using a total number of samples
$T\geq\frac{2SA\gamma^{2}}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(\frac{e^{|\beta|/(1-\gamma)}-1}{|\beta|}\bigg)^{2}\log\bigg(\frac{2SA}{\delta}\bigg)\,.$
The case of $\beta<0$ is proven using very similar lines, but will use inequality (ii) in Lemma 6.1. ∎
6.1.1 Proof of Theorem 2